# Effects of intertube dipole-dipole interactions in nearly integrable one-dimensional 162Dy gases

### v1

@@ Abstract
We study the effects of the intertube dipole-dipole interactions (DDI) in recent experiments with arrays of nearly integrable one-dimensional (1D) dipolar Bose gases of 162Dy atoms. An earlier theoretical modeling ignored those interactions, which we include here via a modification of the 1D confining potentials. We investigate the effects of the intertube DDI both during the state preparation and during the measurements of the rapidity distributions. We explore how the strength of the contact interactions and the magnetic field angles modify the intertube DDI corrections. We find that those corrections slightly change both the properties of the equilibrium state and the rapidity measurements. Remarkably, however, the changes nearly cancel each other, resulting in measured rapidity distributions that are very close to those predicted in the absence of the intertube DDI.
@@ I Introduction
Unlike for generic quantum many-body systems, near-integrability makes possible the existence of long-lived quasiparticles that emerge from interactions among particles. The distribution of the rapidities (the momenta of the quasiparticles), fully characterizes the equilibrium states of nearly integrable systems [1]. Understanding the out-of-equilibrium dynamics of those systems is one of the current research frontiers in quantum many-body physics. Interesting behaviors have been explored, such as: generalized thermalization beyond traditional statistical mechanics paradigms [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], fast prethermal dynamics and slow thermalization [14, 15, 16, 17, 18, 19, 20], generalized hydrodynamics (GHD) [21, 22, 23, 24, 25, 26, 27, 28], and anomalous transport [29, 30, 31, 32, 33]. Well-defined quasiparticles and their rapidity distributions lie at the heart of the theoretical understanding of those behaviors.
@@
Nearly integrable systems have generated much theoretical interest because of their elegant mathematical description, and they have been realized in state-of-the-art ultracold atom experimental platforms [34, 35]. A well-studied example is arrays of 1D bosonic gases with contact interactions trapped in deep two-dimensional (2D) optical lattices [36, 37]. These systems are described by the Lieb-Liniger model [38, 39], which is a paradigmatic integrable model. A recent experimental breakthrough employing a modified time-of-flight (TOF) expansion sequence made possible the measurement of the rapidity distributions in such arrays [40], enabling the observation of the dynamical fermionization of the momentum distribution of Tonks-Girardeau gases [41]. Using this protocol, experiments have tested the applicability of GHD [24], measured the rapidity distributions of dipolar 1D ground-state gases [42], characterized the nonlinear response of quantum many-body scar states [27], studied 1D anyonic correlations [43], and found evidence for Bethe strings [44]. Probing the local rapidity distribution has also been achieved in atom-chip experiments by selectively expanding a portion of the 1D gas [45].
@@
In this work, we focus on the theoretical description of recent experiments with an array of nearly integrable 1D dipolar Bose gases of 162Dy atoms [42]. The long-range dipole-dipole interaction (DDI) between magnetic atoms introduces new features that are not available in 1D Bose gases interacting with purely contact interactions. For example, it provides a tunable integrability breaking term, enabling the study of thermalization near integrability [15, 46]. In addition, a repulsive DDI can stabilize gases with attractive contact interactions and enable the preparation of metastable super-Tonks-Girardeau gas states, including those within the quantum many-body scar regime [47, 48, 27]. The rapidity and momentum distributions of the equilibrium states in the regime with repulsive contact interactions were studied in Ref. [42]. Notably, the theoretical model was able to describe the main experimental observations by accounting for only the intratube DDI (the DDI between bosons in the same 1D gas) via a modification of the contact interaction. The intertube DDI was ignored. (This is the DDI between bosons in different 1D gases confined in nearby tubes of the 2D optical lattice potential.)
@@
In the experimental work of Ref. [42], the first step in the equilibrium state preparation was the loading of the BEC into a deep 2D optical lattice; this described in detail in Sec. II.1. That step was carried out using a magnetic field orientated at an angle at which the intratube DDIs vanish ($\sim$55 with respect to the 1D tube axis). The intertube DDIs, which were present during that experimental step, were neglected in the theoretical modeling. The main goal of this work is to incorporate the intertube DDI into the theoretical modeling of that step in the experiment, and then to account for it in the other steps, including the expansion dynamics used to access the rapidity distributions.
@@
We treat the leading effect of the intertube DDI as a modification to the 1D trapping potentials. Its effect depends on the atom distribution across the entire 2D array of 1D gases. During the equilibrium state preparation, the intertube DDI affects the atom number and the temperature of each 1D gas. Later, it also affects the dynamics of the 1D gases during the expansion used to measure the rapidity distribution. For sufficiently long expansion times, the intertube DDI energy is converted into kinetic energy, and the rapidity distribution is slightly modified from that in the equilibrium state. Remarkably, we find that the changes in the rapidity distribution introduced by the intertube DDI (to the equilibrium state and during the expansion dynamics) nearly cancel each other, resulting in a measured rapidity distribution that is very close to that predicted in the absence of the intertube DDI. Therefore, the DDI is unlikely to be the cause for the disagreement between the model and experimental results in Ref. [42]. Nonthermal effects, due to the near-integrability of the 1D gases, are the most probable cause. We will explore them in future studies.
@@
The manuscript is organized as follows. In Sec. II, we review the experimental setup and the theoretical modeling used in Ref. [42]. We then discuss how we account for the leading effect of the intertube DDI via a modification of the 1D trapping potentials. In Sec. III, we systematically study the effect of the intertube DDI during the equilibrium state preparation in the experiments. In Sec. IV, we discuss how the intertube DDI modifies the (otherwise conserved) rapidity distributions during the 1D expansion. Our main findings are summarized in Sec. VI.
@@ II Experimental setup and its modeling
We review the experimental setup used in Ref. [42] to create arrays of nearly integrable 1D dipolar Bose gases of 162Dy atoms, and how the rapidity distribution measurements were carried out. We then discuss the theoretical modeling used in Ref. [42], and how in this work we calculate the leading corrections introduced by the intertube DDI.
@@ II.1 Experimental sequence
Figure 1: (a) Spatial depiction of the dipolar 1D gases. The 2D array (in the $y$-$z$ plane) of 1D gases (along the $x$ direction) is created using a deep 2D optical lattice with wavelength $\lambda$. The atomic dipoles align with the magnetic field $\bf B$ producing the intertube and intratube DDIs. (b) Sequence used to measure the rapidity distributions. After preparing the equilibrium 1D gases (tubes), the trapping potential $V_{1D}$ is turned off and the atoms are allowed to expand in one dimension in the presence of interactions. The momentum distribution of the 1D gases becomes equal to the rapidity distribution after a long-enough expansion time $t_{\text{ev}}$. (c) Theoretical modeling of the experimentally created equilibrium state. Starting from a BEC (left), the 2D optical lattice potential $U_{2D}$ is ramped up to create the array of tubes (right). We assume that at $U_{2D}^{*}$ all the tubes decouple from each other, and that they are all at the same temperature $T_{2D}^{*}$ (center). The ramping of 2D lattice is assumed to be adiabatic for $U_{2D}>U_{2D}^{*}$.
The rapidity measurement is performed by turning off the 1D trapping potential $V_{1D}$ while keeping the 2D lattice $U_{2D}$ on and allowing the 1D gases to expand along the 1D direction for a time $t_{\text{ev}}$; see Fig. 1(b). After a long expansion time $t_{\text{ev}}$, the momentum distribution of the 1D gases approaches the rapidity distribution [41, 49, 50, 51, 52, 53]. A standard 3D TOF momentum measurement is then used to measure the asymptotic momentum distribution of all the atoms and therefore the rapidity distribution.
@@ II.2 Modeling uncoupled 1D gases
The theoretical model used in Ref. [42] neglected the intertube DDI; i.e., it treated the experimental system as a 2D array of independent tubes. Within this approximation, each 1D gas can be described using the extended Lieb-Liniger model [38],
$H$ $=\sum_{j=1}^{N}\left[-\frac{\hbar^{2}}{2\text{m}}\frac{\partial^{2}}{\partial x_{j}^{2}}+V_{\rm 1D}(x_{j})\right]$
@@
which accounts for the confining potential $V_{\rm 1D}$ along the 1D direction and the intratube DDI $U^{\rm intra}_{\rm DDI}$. The effective 1D contact interaction $g_{\rm 1D}^{\rm vdW}$ is the result of the van der Waals force and depends on the depth of the 2D optical lattice $U_{\text{2D}}$ and the magnetic filed (which determines the 3D $s$-wave scattering length $a_{\rm 3D}$). The explicit form of $g_{\rm 1D}^{\rm vdW}$ is
$g_{\rm 1D}^{\rm vdW}=-\frac{\hbar^{2}}{\text{m}a_{\rm 1D}}\,$, (2)
where $a_{\rm 1D}=-a_{\perp}(a_{\perp}/a_{\rm 3D}-C)/2$ is the 1D scattering length, $a_{\perp}=\sqrt{2\hbar/(\text{m}\omega_{\perp})}$ is the transverse confinement, $\omega_{\perp}=\sqrt{2U_{\rm 2D}k_{0}^{2}/\text{m}}$, and $C=-\zeta(1/2)\simeq 1.4603$ [39].
@@
$U^{\rm intra}_{\rm DDI}(\theta_{\text{B}},x_{j}-x_{l})=\frac{\mu_{0}\mu^{2}}{4\pi}\frac{1-3\cos^{2}\theta_{\text{B}}}{\sqrt{2}a_{\perp}^{3}}\bigg[V^{\rm intra}_{\rm DDI}(u)-\frac{8}{3}\delta(u)\bigg]\,$, (3)
where $V^{\rm intra}_{\rm DDI}(u)=-2|u|+\sqrt{2\pi}(1+u^{2})e^{u^{2}/2}{\rm erfc}(|u|/\sqrt{2})$, $u=\sqrt{2}(x_{j}-x_{l})/a_{\perp}$, and ${\rm erfc}(x)$ is the complementary error function [54, 55, 15, 56, 57]. The dipole moment of 162Dy atom is $\mu=9.93\mu_{\text{B}}$. In Ref. [42], $\theta_{B}=55^{\circ}$ during the ramping up of the 2D lattice, so that $U^{\rm intra}_{\rm DDI}\approx 0$.
To account for the leading (short-range) effect of the intratube DDI, we treat it as a modification to the contact interaction [42, 15, 27],
$\tilde{U}^{\rm intra}_{\text{DDI}}(\theta_{\text{B}},x_{j}-x_{l})$ $=\frac{\mu_{0}\mu^{2}}{4\pi}\frac{1-3\cos^{2}\theta_{\text{B}}}{2a_{\perp}^{2}}\left[A-\frac{8}{3}\right]\delta(x_{j}-x_{l})$
@@
$\tilde{H}=\sum_{j=1}^{N}\bigg[-\frac{\hbar^{2}}{2\text{m}}\frac{\partial^{2}}{\partial x_{j}^{2}}+V_{\rm 1D}(x_{j})\bigg]+\sum_{1\leq j<l\leq N}g_{\rm 1D}\delta(x_{j}-x_{l})\,$, (5)
where the effective 1D contact interaction is $g_{\rm 1D}=g_{\rm 1D}^{\rm vdW}+g_{\rm 1D}^{\rm DDI}$. In the absence of $V_{\rm 1D}$, Hamiltonian (5) is integrable and it is exactly solvable via the Bethe ansatz [38, 58]. The ground-state observables depend only on the dimensionless parameter $\gamma=\text{m}g_{\text{1D}}/\hbar^{2}n_{\text{1D}}$, where $n_{\text{1D}}$ is the 1D density. When $V_{\rm 1D}$ is present, density and rapidity distributions can be obtained using a local density approximation (LDA).
@@ II.3 Modeling the state preparation in uncoupled 1D gases
To describe the experimental results, in Ref. [42] we averaged the observables over all the tubes. We assumed that each tube is in thermal equilibrium after the state preparation, so the parameters that need to be determined, for each tube $l$ at position $(y_{l},z_{l})$, are the number of atoms $N_{l}$ and the temperature $T_{l}$. In Fig. 1(c), we illustrate the modeling and assumptions used to determine $N_{l}$ and $T_{l}$.
Before turning on the 2D optical lattice, the system is in a 3D BEC with a total atom number $N_{\rm tot}$ and a temperature $T_{\rm 3D}$ (left panel). As the 2D lattice is ramped up, the atoms begin to be confined in coupled (number fluctuating) quasi-1D gases. We assumed that, at a certain lattice depth $U^{*}_{\rm 2D}$, the system decouples into independent 1D gases with a fixed number of atoms (middle panel). We also assumed that all the tubes are in thermal equilibrium at the same temperature $T^{*}_{\rm 2D}$ at decoupling.
Under these assumptions, and having knowledge of $N_{\rm tot}$ and the confining potentials along all directions, we used the thermodynamic Bethe ansatz [58] (TBA, which describes 1D homogeneous Bose gases with contact interactions at finite temperature) and the LDA (to account for the confining potentials) to determine the atom number $N_{l}$ and entropy $S^{*}_{l}$ in each tube at the decoupling point for a dense grid of values of $U^{*}_{\rm 2D}$ and $T^{*}_{\rm 2D}$. $U^{*}_{\rm 2D}$ determines the contact interaction parameter $g^{*}_{1D}$ via Eq. (2). We stress that, in the experiments, it is unlikely that all the tubes decouple at a single lattice depth $U^{*}_{\rm 2D}$. However, in Ref. [42], we found that the theoretical results are not very sensitive to the decoupling depth $U^{*}_{\rm 2D}$ near to the optimal $U^{*}_{\rm 2D}$ and $T^{*}_{\rm 2D}$. (However, they are more sensitive to the value of $T^{*}_{\rm 2D}$.) Therefore, the assumption of a single decoupling depth is expected to be a reasonable approximation for our system.
Next, we sequentially determined the temperatures of the 1D gases at the end of the optical lattice ramp [Fig. 1(c), right panel], after reducing the trapping frequency ($f_{x}$) and after the magnetic field has been adjusted. To do so, we also use the TBA, LDA and assume that all the changes made after the decoupling depth $U^{*}_{\rm 2D}$ are adiabatic (i.e., that the entropy $S_{l}$ of each tube does not change) and that the 1D gases remain in thermal equilibrium throughout. Using the final temperatures and model parameters after all the experimental changes, we calculated the experimental observables. The two free parameters in our model, $U^{*}_{2D}$ and $T^{*}_{2D}$, were fixed by minimizing the difference between theoretical and experimental momentum and rapidity distributions.

### v2

@@ Abstract
We study the effects of the intertube dipole-dipole interactions (DDI) in recent experiments with arrays of nearly integrable one-dimensional (1D) dipolar Bose gases of 162Dy atoms. An earlier theoretical modeling ignored those interactions, which we include here via a modification of the 1D confining potentials. We investigate the effects of the intertube DDI both during the state preparation and during the measurements of the rapidity distributions. We explore how the strength of the contact interactions and the magnetic-field angles modify the intertube DDI corrections. We find that those corrections slightly change both the properties of the equilibrium state and the rapidity measurements. Remarkably, however, the changes nearly cancel each other, resulting in measured rapidity distributions that are very close to those predicted in the absence of the intertube DDI.
@@ I Introduction
Unlike for generic quantum many-body systems, near-integrability makes possible the existence of long-lived quasiparticles that emerge from interactions among particles. The distribution of the rapidities (the momenta of the quasiparticles) fully characterizes the equilibrium states of nearly integrable systems [1]. Understanding the out-of-equilibrium dynamics of those systems is one of the current research frontiers in quantum many-body physics. Interesting behaviors have been explored such as generalized thermalization beyond traditional statistical mechanics paradigms [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], fast prethermal dynamics and slow thermalization [14, 15, 16, 17, 18, 19, 20], generalized hydrodynamics (GHD) [21, 22, 23, 24, 25, 26, 27, 28, 29], and anomalous transport [30, 31, 32, 33, 34]. Well-defined quasiparticles and their rapidity distributions lie at the heart of the theoretical understanding of those behaviors.
@@
Nearly integrable systems have generated much theoretical interest because of their elegant mathematical description and they have been realized in state-of-the-art ultracold-atom experimental platforms [35, 36]. A well-studied example is arrays of one-dimensional (1D) bosonic gases with contact interactions trapped in deep 2D optical lattices [37, 38]. These systems are described by the Lieb-Liniger model [39, 40], which is a paradigmatic integrable model. A recent experimental breakthrough employing a modified time-of-flight (TOF) expansion sequence made possible the measurement of the rapidity distributions in such arrays [41], enabling the observation of the dynamical fermionization of the momentum distribution of Tonks-Girardeau gases [42]. Using this protocol, experiments have tested the applicability of GHD [24], measured the rapidity distributions of dipolar 1D ground-state gases [43], characterized the nonlinear response of quantum many-body scar states [27], studied 1D anyonic correlations [44], and found evidence for Bethe strings [29]. Probing the local rapidity distribution has also been achieved in atom-chip experiments by selectively expanding a portion of the 1D gas [45].
@@
In this work we focus on the theoretical description of recent experiments with an array of nearly integrable 1D dipolar Bose gases of 162Dy atoms [43]. The long-range dipole-dipole interaction (DDI) between magnetic atoms introduces new features that are not available in 1D Bose gases interacting with purely contact interactions. For example, it provides a tunable integrability breaking term, enabling the study of thermalization near integrability [15, 46]. In addition, a repulsive DDI can stabilize gases with attractive contact interactions and enable the preparation of metastable super-Tonks-Girardeau gas states, including those within the quantum many-body scar regime [47, 48, 27]. The rapidity and momentum distributions of the equilibrium states in the regime with repulsive contact interactions were studied in Ref. [43]. Notably, the theoretical model was able to describe the main experimental observations by accounting for only the intratube DDI (the DDI between bosons in the same 1D gas) via a modification of the contact interaction. The intertube DDI was ignored. (This is the DDI between bosons in different 1D gases.)
@@
In the experimental work of Ref. [43], the first step in the equilibrium state preparation was loading the Bose-Einstein condensate (BEC) into a deep 2D optical lattice (see Sec. II.1). That step was carried out using a magnetic field oriented at an angle at which the intratube DDIs vanish (approximately 55 with respect to the 1D tube axis). The intertube DDIs, which were present during that experimental step, were neglected in the theoretical modeling. The main goal of this work is to incorporate the intertube DDI into the theoretical modeling of that step in the experiment and then to account for it in the other steps, including the expansion used to access the rapidity distributions. While we expect the effect of the intertube DDI to be small in each stage of the modeling, it is important to understand how those effects add up.
@@
We treat the leading effect of the intertube DDI as a modification to the 1D trapping potentials. It depends on the atom distribution across the entire 2D array of 1D gases. During the equilibrium state preparation, the intertube DDI affects the atom number and the temperature of each 1D gas. Later, it also affects the dynamics of the 1D gases during the expansion used to measure the rapidity distribution. For sufficiently long expansion times, the intertube DDI energy is converted into kinetic energy, and the rapidity distribution is slightly modified from that in the equilibrium state. Remarkably, we find that the changes in the rapidity distribution introduced by the intertube DDI (to the equilibrium state and during the expansion dynamics) nearly cancel each other, resulting in a measured rapidity distribution that is very close to that predicted in the absence of the intertube DDI. Therefore, the DDI is unlikely to be the cause of the disagreement between the model and experimental results in Ref. [43]. Nonthermal effects, due to the near-integrability of the 1D gases, are the most probable cause; we leave this exploration to future studies.
@@
The paper is organized as follows. In Sec. II we review the experimental setup and the theoretical modeling used in Ref. [43]. We then discuss how we account for the leading effect of the intertube DDI via a modification of the 1D trapping potentials. In Sec. III we systematically study the effect of the intertube DDI during the equilibrium state preparation in the experiments. In Sec. IV we discuss how the intertube DDI modifies the (otherwise conserved) rapidity distributions during the 1D expansion. Our main findings are summarized in Sec. VI.
@@ II Experimental setup and its modeling
We review the experimental setup used in Ref. [43] to create arrays of nearly integrable 1D dipolar Bose gases of 162Dy atoms and how the rapidity distribution measurements were carried out. We then discuss the theoretical modeling used in Ref. [43] and how in this work we calculate the leading corrections introduced by the intertube DDI.
@@ II.1 Experimental sequence
Figure 1: (a) Spatial depiction of the dipolar 1D gases. The 2D array (in the $y$-$z$ plane) of 1D gases (in the $x$ direction) is created using a deep 2D optical lattice with wavelength $\lambda$. The atomic dipoles align with the magnetic field $\bf B$ producing the intertube and intratube DDIs. (b) Sequence used to measure the rapidity distributions. After preparing the equilibrium 1D gases (tubes), the trapping potential $V_{\mathrm{1D}}$ is turned off and the atoms are allowed to expand in one dimension in the presence of interactions. The momentum distribution of the 1D gases becomes equal to the rapidity distribution after a long-enough expansion time $t_{\text{ev}}$. (c) Theoretical modeling of the experimentally created equilibrium state. Starting from a BEC (left), the 2D optical lattice potential $U_{\mathrm{2D}}$ is ramped up to create the array of tubes (right). We assume that at $U_{\mathrm{2D}}^{*}$ all the tubes decouple from each other, and that they are all at the same temperature $T_{\mathrm{2D}}^{*}$ (center). The ramping of 2D lattice is assumed to be adiabatic for $U_{\mathrm{2D}}>U_{\mathrm{2D}}^{*}$.
The rapidity measurement is done turning off the 1D trapping potential $V_{\mathrm{1D}}$ while keeping the 2D lattice $U_{2D}$ on and allowing the 1D gases to expand in the 1D direction for a time $t_{\text{ev}}$ [see Fig. 1(b)]. After a long expansion time $t_{\text{ev}}$, the momentum distribution of the 1D gases approaches the rapidity distribution [42, 49, 50, 51, 52, 53]. A 3D TOF momentum measurement is then used to measure the asymptotic momentum distribution of all the atoms and therefore the rapidity distribution.
@@ II.2 Modeling uncoupled 1D gases
The theoretical model used in Ref. [43] neglected the intertube DDI, i.e., it treated the experimental system as a 2D array of independent tubes. Within this approximation, each 1D gas can be described using the extended Lieb-Liniger model [39],
$H$ $=\sum_{j=1}^{N}\left(-\frac{\hbar^{2}}{2m}\frac{\partial^{2}}{\partial x_{j}^{2}}+V_{\rm 1D}(x_{j})\right)$
@@
which accounts for the confining potential $V_{\rm 1D}$ in the 1D direction and the intratube DDI $U^{\rm intra}_{\rm DDI}$ between the $N$ atoms. The effective 1D contact interaction $g_{\rm 1D}^{\rm vdW}$ is the result of the van der Waals force and depends on the depth of the 2D optical lattice $U_{\text{2D}}$ and the magnetic field (which determines the 3D $s$-wave scattering length $a_{\rm 3D}$). The explicit form of $g_{\rm 1D}^{\rm vdW}$ is
$g_{\rm 1D}^{\rm vdW}=-\frac{\hbar^{2}}{ma_{\rm 1D}}\,$, (2)
where $a_{\rm 1D}=-a_{\perp}(a_{\perp}/a_{\rm 3D}-C)/2$ is the 1D scattering length, $a_{\perp}=\sqrt{2\hbar/(m\omega_{\perp})}$ is the transverse confinement, $\omega_{\perp}=\sqrt{2U_{\rm 2D}k_{0}^{2}/m}$, and $C=-\zeta(1/2)\simeq 1.4603$ [40].
@@
$U^{\rm intra}_{\rm DDI}(\theta_{\text{B}},x_{j}-x_{l})=\frac{\mu_{0}\mu^{2}}{4\pi}\frac{1-3\cos^{2}\theta_{\text{B}}}{\sqrt{2}a_{\perp}^{3}}\bigg(V^{\rm intra}_{\rm DDI}(u)-\frac{8}{3}\delta(u)\bigg)\,$, (3)
where $V^{\rm intra}_{\rm DDI}(u)=-2|u|+\sqrt{2\pi}(1+u^{2})e^{\frac{u^{2}}{2}}{\rm erfc}(\tfrac{|u|}{\sqrt{2}})$, $u=\tfrac{\sqrt{2}}{a_{\perp}}(x_{j}-x_{l})$, ${\rm erfc}(x)$ is the complementary error function [54, 55, 15, 47, 56], and $\mu=9.93\mu_{\text{B}}$ is the 162Dy atom dipole moment.
In Ref. [43], the 2D optical lattice used to created the array of 1D gases was ramped up such that $\theta_{B}=55^{\circ}$, which means that $U^{\rm intra}_{\rm DDI}\approx 0$ during this process (studied in Sec. III) and in the ensuing measurements of the rapidity distribution (studied in Sec. IV). Also, $U^{\rm intra}_{\rm DDI}\approx 0$ for initial states and their corresponding rapidity measurements in which different values of $g_{\rm 1D}^{\rm vdW}$ were explored experimentally using a Feshbach resonance at fixed $\theta_{B}=55^{\circ}$ (studied in Sec. V.1).
To account for the leading (short-range) effect of the intratube DDI when $\theta_{B}\neq 55^{\circ}$ (studied in Sec. V.2), we treat it as a modification to the contact interaction [43, 15, 27],
$\tilde{U}^{\rm intra}_{\text{DDI}}(\theta_{B},x_{j}-x_{l})$ $=\frac{\mu_{0}\mu^{2}}{4\pi}\frac{1\!-\!3\cos^{2}\theta_{B}}{2a_{\perp}^{2}}\!\left(\!A-\frac{8}{3}\right)\!\delta(x_{j}\!-\!x_{l})$
@@
$\tilde{H}=\sum_{j=1}^{N}\bigg(-\frac{\hbar^{2}}{2m}\frac{\partial^{2}}{\partial x_{j}^{2}}+V_{\rm 1D}(x_{j})\bigg)+\sum_{1\leq j<l\leq N}g_{\rm 1D}\delta(x_{j}-x_{l})\,$, (5)
where the effective 1D contact interaction is $g_{\rm 1D}=g_{\rm 1D}^{\rm vdW}+g_{\rm 1D}^{\rm DDI}$. In the absence of $V_{\rm 1D}$, Hamiltonian (5) is integrable and it is exactly solvable via the Bethe ansatz [39, 57]. The ground-state observables depend only on the dimensionless parameter $\gamma=mg_{\text{1D}}/(\hbar^{2}n_{\text{1D}})$, where $n_{\text{1D}}$ is the 1D density. When $V_{\rm 1D}$ is present, density and rapidity distributions can be obtained using a local density approximation (LDA).
@@ II.3 Modeling the state preparation in uncoupled 1D gases
To describe the experimental results, in Ref. [43] we averaged the observables over all the tubes. We assumed that each tube is in thermal equilibrium after the state preparation, so the parameters that need to be determined, for each tube $\ell$ at position $(y_{\ell},z_{\ell})$, are the number of atoms $N_{\ell}$ and the temperature $T_{\ell}$. In Fig. 1(c) we illustrate the modeling and assumptions used to determine $N_{\ell}$ and $T_{\ell}$.
Initially, the system is in a 3D BEC with total atom number $N_{\rm tot}$ and temperature $T_{\rm 3D}$ (left panel). As the 2D lattice is ramped up, the atoms are confined in coupled (number fluctuating) quasi-1D gases. We assumed that, at a lattice depth $U^{*}_{\rm 2D}$, the system decouples into independent 1D gases with fixed atom number (middle panel), and that all the tubes are in thermal equilibrium at the same temperature $T^{*}_{\rm 2D}$.
Under these assumptions and having knowledge of $N_{\rm tot}$ and the confining potentials in all directions, we used the thermodynamic Bethe ansatz [57] (TBA) (which describes 1D homogeneous Bose gases with contact interactions at finite temperature) and the LDA (to account for the confining potentials) to determine the atom number $N_{\ell}$ and entropy $S^{*}_{\ell}$ in each tube at the decoupling point for a dense grid of values of $U^{*}_{\rm 2D}$ and $T^{*}_{\rm 2D}$. The lattice depth $U^{*}_{\rm 2D}$ determines the contact interaction parameter $g^{*}_{\rm 1D}$ via Eq. (2). We stress that, in the experiments, it is unlikely that all the tubes decouple at a single lattice depth $U^{*}_{\rm 2D}$. However, in Ref. [43] the theoretical results were found to be robust to changes in decoupling depth $U^{*}_{\rm 2D}$ near the optimal $U^{*}_{\rm 2D}$ and $T^{*}_{\rm 2D}$. (They are more sensitive to the value of $T^{*}_{\rm 2D}$.) Therefore, the assumption of a single decoupling depth is expected to be a reasonable approximation for our system.
Next we sequentially determined the temperatures of the 1D gases at the end of the optical lattice ramp [Fig. 1(c), right panel], after reducing the trapping frequency $f_{x}$ and after the magnetic field has been adjusted. To do so, we also use the TBA and LDA and assume that all the changes made after the decoupling depth $U^{*}_{\rm 2D}$ are adiabatic (i.e., that the entropy $S_{\ell}$ of each tube does not change) and that the 1D gases remain in thermal equilibrium throughout. Using the final temperatures and model parameters after all the experimental changes, we calculated the experimental observables. The two free parameters in our model, $U^{*}_{\rm 2D}$ and $T^{*}_{\rm 2D}$, were fixed by minimizing the difference between theoretical and experimental momentum and rapidity distributions.